# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

### v3

@@ 4.2 Mass and stiffness matrix structure: estimates on $\tau_{K}$
Suppose $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are of SPD-type and $\mathbb{D}_{K}$ satisfies ANAC. If ${\boldsymbol{b}}$ has nonpositive divergence, then $A_{ii}$ and $M_{ii}$ are both strictly positive. If ${\boldsymbol{b}}$ has strictly positive divergence, then $A_{ii}>0$ if
$\tau_{K}>-\frac{\varphi_{i}}{{\boldsymbol{b}}\cdot\nabla\varphi_{i}}$ (79)
and $M_{ii}>0$ if
$\tau_{K}<\frac{4|K|}{(d+1)(d+2)}\left(\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}\right)^{-1}.$ (80)
@@
We consider a patch of elements $\Omega_{i}$ that share a vertex ${\boldsymbol{a}}_{i}$. Recall the definition of $A_{ii}$ we have
$A_{ii}$ $=\sum_{K\in\Omega_{i}}\int_{K}\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{i}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{i}})^{2}+\tau_{K}\left(\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}\right)\cdot\nabla({\boldsymbol{b}\cdot\nabla\varphi_{i})}\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}$ (81)
We first concentrate on the first two integrals and make use of $\tau_{K}$ in order to bound these integrals below by zero. Note that by (43) the first integral is nonnegative if ${\boldsymbol{b}}$ has nonpositive divergence. Then
$\int_{K}(\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})^{2})\textrm{d}{\boldsymbol{x}}\geq 0.$ (82)
If ${\boldsymbol{b}}$ has strictly positive divergence, then (82) holds under the condition (79). Indeed, (43) implies
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}<0$ (83)
and hence ${\boldsymbol{b}}\cdot\nabla\varphi_{i}<0$ since $\varphi_{i}>0$ on element $K\in\Omega_{i}$ and $\tau_{K}>0$ for all $K\in\Omega$. By factoring out the convection term, the lower bound on $\tau_{K}$ given in equation (79) gives (82).
We now turn our attention to the third term in (81). Let $\lambda_{j}$ and ${\boldsymbol{\omega}}_{j}$ be the quadrature weights and nodes, respectively on element $K\in\Omega_{i}$. Then
$\int_{K}(\mathbb{D}\nabla\varphi_{i})\cdot\nabla({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}$ $\approx|K|\sum_{i=1}^{m}\lambda_{j}(\nabla({\boldsymbol{b}}({\boldsymbol{\omega}}_{j})\cdot\nabla\varphi_{i})^{T}\mathbb{D}({\boldsymbol{\omega}}_{j})\nabla\varphi_{i}$ (84)
$=|K|(\nabla\varphi_{i})^{T}\mathbb{K}_{K}(\nabla\varphi_{i})$ (85)
It follows from Lemma 4.2 that $A_{ii}>0$.
On the other hand,
$M_{ii}$ $=\sum_{K\in\Omega_{i}}(\varphi_{i},\varphi_{i}+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i}))_{0,K}$ (86)
$=\sum_{K\in\Omega_{i}}\int_{K}\left(\varphi_{i}^{2}+\tau_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\right)\textrm{d}{\boldsymbol{x}}$ (87)
$=\sum_{K\in\Omega_{i}}\left(\frac{2|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right).$ (88)
As we saw previously, the integral is nonnegative if ${\boldsymbol{b}}$ has nonpositive divergence. Otherwise, we apply the upper bound on $\tau_{K}$ given by (80). Hence, $M_{ii}>0$. ∎
      Remark 4.2.
To summarize the previous lemma, under the ANAC and SPD-type conditions, if ${\boldsymbol{b}}$ is nonpositive, then $A_{ii}$ and $M_{ii}>0$. Otherwise, we require
$-\frac{\varphi_{i}}{{\boldsymbol{b}\cdot\nabla\varphi_{i}}}<\tau_{K}<\frac{4|K|}{(d+1)(d+2)}\left(\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}\right)^{-1}.$ (89)
@@
If $\mathbb{K}_{K}$ and ${\mathbb{D}}_{K}$ are of SPD-type, satisfy ANAC, and if the following two conditions are met
$\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}^{2}}{\rho}\leq\cos(\check{\alpha}_{ij})$ (90)
@@
$\frac{2h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}}{\rho\sqrt{(d+1)(d+2)}}\leq\cos(\tilde{\alpha}_{ij}),$ (91)
where $\rho=\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}$, then $A_{ij}\leq 0$, for $i\neq j$.
Proof.
For the ${\boldsymbol{A}}$ matrix, we have
$A_{ij}=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\int_{K}\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{j}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{j}})({\boldsymbol{b}\cdot\nabla\varphi_{i}})$ $+\tau_{K}\left(\varepsilon\mathbb{D}_{K}\nabla\varphi_{j}\right)\cdot\nabla({\boldsymbol{b}\cdot\nabla\varphi_{i})}\textrm{d}{\boldsymbol{x}}$ (92)
$+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{j}$ (93)
We investigate the three terms in the integral of $A_{ij}$. The first integral can be simplified using Cauchy-Schwarz and Holder’s inequalities along with (42),
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{j})\textrm{d}{\boldsymbol{x}}$ $\leq\sqrt{\frac{|K|}{(d+1)(d+2)}}\|{\boldsymbol{b}}\|_{\infty,K}\|\nabla\varphi_{i}\|_{2,K}$ (94)
$=\sqrt{\frac{2}{(d+1)(d+2)}}\left(\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right).$ (95)
For the second term, we again use Cauchy-Schwarz and Holder’s inequality and have the following bound on the convection integral
$\int_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})({\boldsymbol{b}}\cdot\nabla\varphi_{j})\textrm{d}{\boldsymbol{x}}$ $\leq\|{\boldsymbol{b}}\|_{\infty,K}^{2}\frac{|K|}{h_{i}^{K}h_{j}^{K}}.$ (96)
We have
$A_{ij}$ $\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\left(\frac{\sqrt{2}|K|}{h_{j}^{K}\sqrt{(d+1)(d+2)}}\|{\boldsymbol{b}}\|_{\infty,K}+\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}^{2}\frac{|K|}{h_{i}^{K}h_{j}^{K}}-\tau_{K}\frac{\cos(\check{\alpha}_{ij}^{K})}{\check{h}_{i}^{K}\check{h}_{j}^{K}}-|K|\frac{\cos(\tilde{\alpha}_{ij}^{K})}{\tilde{h}_{i}^{K}\tilde{h}_{j}^{K}}\right)$ (97)
$\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{\sqrt{2}}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)$ (98)
$\hskip 144.54pt-\frac{1}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij}^{K})\rho_{min}(\mathbb{K}_{K})+|K|\cos(\tilde{\alpha}_{ij}^{K})\rho_{min}(\varepsilon\mathbb{D}_{K})\right)$ (99)
$\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{\sqrt{2}}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)$ (100)
$\hskip 144.54pt-\frac{\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)$ (101)
By (90) and (91), $A_{ij}\leq 0$. ∎
@@
If
$\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\geq 0$ (102)
or
$\tau_{K}\leq-\frac{|K|}{\displaystyle(d+1)(d+2)}\left(\int_{K}\varphi_{j}(\boldsymbol{b}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right)^{-1},$ (103)
@@
The is immediate based on the definition of $M_{ij}$ and (21). ∎
Combining the previous two lemmas, we have the following result.
@@
Let $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ be of SPD-type, satisfy ANAC, and (90) and (91) be satisfied. The following nonnegative condition
$M_{ij}-\Delta t(1-\theta)A_{ij}\geq 0$ (104)
holds for the off-diagonal entries if the convection integral is nonnegative or $\tau_{K}$ satisfies (103).
@@ 4.3 Temporal estimates
Suppose $\Delta t$ satisfies
$\Delta t\geq-\frac{1}{\theta\sqrt{(d+1)(d+2)}}\max_{K\in\mathcal{T}_{h}}\max_{\underset{i\neq j}{i,j=1,\dots,d+1}}\left(\frac{|K|h_{j}^{K}\left(h_{i}^{K}+2\tau_{K}\sqrt{(d+1)(d+2)}\|{\boldsymbol{b}}\|_{\infty,K}\right)}{2|K|h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}+\tau_{K}h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}^{2}-\rho(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij}))}\right),$ (105)
where $\rho=\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}$. Suppose also that $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are of SPD-type, satisfy ANAC, and satisfy (90) and (91) (without equality). If, for each element $K$, ${\boldsymbol{b}}$ has nonpositive divergence on or $\tau_{K}$ satisfies (89), then $\tilde{\boldsymbol{A}}$ is an $M$-matrix.
@@
From Lemma 4.4, the diagonal entries of ${\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}}$ is strictly positive. We show
$M_{ij}+\theta\Delta tA_{ij}\leq 0,\ \ \forall i\neq j,i=1,\dots,N_{vi},j=1,\dots,N_{v},$ (106)
since this is equivalent to showing that $M+\theta\Delta tA$ is a $Z$-matrix and an $M$-matrix is a $Z$-matrix that is inverse positive.
For the ${\boldsymbol{M}}$ matrix, we have

### v4

@@ 4.2 Mass and stiffness matrix structure: estimates on $\tau_{K}$
Suppose $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are symmetric and positive definite and the meshes associated with $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC. Then the diagonal entries of the mass and stiffness matrices satisfy $M_{ii}>0$ and $A_{ii}>0$ provided
$\tau_{K}\|b\|_{\infty,K}\leq\frac{h_{i}^{K}}{d+2},$ (86)
and
$\frac{\rho_{\min}(\mathbb{D}_{K})+\tau_{K}\rho_{\min}(\mathbb{K}_{K})}{(h_{i}^{K})^{2}}>\frac{1}{2\varepsilon\tau_{K}(d+1)(d+2)}.$ (87)
@@
We consider a patch of elements $\omega_{i}$ that share a vertex ${\boldsymbol{a}}_{i}$. Using (49) and (21), from (26), we obtain
$M_{ii}\geq\sum_{K\in\omega_{i}}\left(\frac{2|K|}{(d+1)(d+2)}-\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\|b\|_{\infty,K}\right),$
which is positive under (86).
Applying Young’s inequality
$\phi_{i}(b\cdot\nabla\phi_{i})+\tau_{K}(b\cdot\nabla\phi_{i})^{2}\geq-\frac{1}{4\tau_{K}}\phi_{i}^{2},$
for $A_{ii}$ in (27), we have,
$A_{ii}$ $=\sum_{K\in\omega_{i}}\int_{K}\left(\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{i}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{i}})^{2}+\tau_{K}\left(\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}\right)\cdot\nabla({\boldsymbol{b}\cdot\nabla\varphi_{i})}\right)\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}$
$=\sum_{K\in\omega_{i}}\int_{K}\left(\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{i}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{i}})^{2}\right)\textrm{d}{\boldsymbol{x}}+\tau_{K}|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{K}_{K}\nabla\varphi_{i}+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}$
$\geq\sum_{K\in\omega_{i}}\int_{K}-\frac{1}{4\tau_{K}}\phi_{i}^{2}\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}\varepsilon(\tau_{K}\mathbb{K}_{K}+\mathbb{D}_{K})\nabla\varphi_{i}$
$=\sum_{K\in\omega_{i}}-\frac{|K|}{2\tau_{K}(d+1)(d+2)}+|K|(\nabla\varphi_{i})^{T}\varepsilon(\tau_{K}\mathbb{K}_{K}+\mathbb{D}_{K})\nabla\varphi_{i}$
$\geq\sum_{K\in\omega_{i}}|K|\left(\varepsilon\frac{\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K})}{(h_{i}^{K})^{2}}-\frac{1}{2\tau_{K}(d+1)(d+2)}\right),$ (88)
which yields positivity under (87). ∎
@@
Suppose $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are symmetric and positive definite and the meshes associated with $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC. For $i\neq j$, define
$\delta_{ij}^{K}=\rho_{\min}(\varepsilon\mathbb{D}_{K})\cos(\widetilde{\alpha}_{ij}^{\mathbb{D}})+\tau_{K}\rho_{\min}(\mathbb{K}_{K})\cos(\check{\alpha}_{ij}^{\mathbb{K}})-\frac{h_{i}^{K}}{d+1}\|b\|_{\infty,K}-\tau_{K}\|b\|_{\infty,K}^{2},$ (89)
where $\widetilde{\alpha}_{ij}^{\mathbb{D}}$ and $\widetilde{\alpha}_{ij}^{\mathbb{K}}$ are the angles in the meshes associated with $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$, respectively. If
$\delta_{ij}^{K}\geq 0\quad\text{ for all }K\in\mathcal{T}_{h},\,i\neq j,$ (90)
then $A_{ij}\leq 0$.
Proof.
For $i\neq j$, from (27) we have
$A_{ij}=\sum_{K\in\omega_{i}\cap\omega_{j}}\left(\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{j})\textrm{d}{\boldsymbol{x}}+\tau_{K}\int_{K}({\boldsymbol{b}\cdot\nabla\varphi_{j}})({\boldsymbol{b}\cdot\nabla\varphi_{i}})\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}(\varepsilon\mathbb{D}_{K}+\varepsilon\tau_{K}\mathbb{K}_{K})\nabla\varphi_{j}\right).$
Applying Lemma 4.2 to $\varepsilon\mathbb{D}_{K}$ and $\varepsilon\mathbb{K}_{K}$, we have
$|K|\nabla\varphi_{i})^{T}(\varepsilon\mathbb{D}_{K})\nabla\varphi_{j}\leq-\frac{|K|}{h_{i}^{K}h_{j}^{K}}\rho_{min}(\varepsilon\mathbb{D}_{K})\cos(\widetilde{\alpha}_{ij}^{\mathbb{D}}),$ (91)
@@
$|K|\nabla\varphi_{i})^{T}(\varepsilon\mathbb{K}_{K})\nabla\varphi_{j}\leq-\frac{|K|}{h_{i}^{K}h_{j}^{K}}\rho_{min}(\varepsilon\mathbb{K}_{K})\cos(\widetilde{\alpha}_{ij}^{\mathbb{K}}).$ (92)
Using inequalities (49) and (51), we have
$A_{ij}$ $\leq\sum_{K\in\omega_{i}\cap\omega_{j}}\frac{|K|}{h_{i}^{K}h_{j}^{K}}\left(\frac{h_{i}^{K}}{d+1}\|b\|_{\infty,K}+\tau_{K}\|b\|^{2}_{\infty,K}-\rho_{min}(\varepsilon\mathbb{D}_{K})\cos(\widetilde{\alpha}_{ij}^{\mathbb{D}})-\tau_{K}\rho_{min}(\varepsilon\mathbb{K}_{K})\cos(\widetilde{\alpha}_{ij}^{\mathbb{K}})\right)$
$=-\sum_{K\in\omega_{i}\cap\omega_{j}}\frac{|K|}{h_{i}^{K}h_{j}^{K}}\delta_{ij}^{K}.$ (93)
Therefore, $A_{ij}\leq 0$ if (90) holds. ∎
In real computations, there is only one mesh at each physical time, thus $\widetilde{\alpha}_{ij}^{\mathbb{D}}$ and $\widetilde{\alpha}_{ij}^{\mathbb{K}}$ will be referred to the same angle, which will be discussed in later subsection 4.4.
@@
For $i\neq j$, if
$\tau_{K}\|b\|_{\infty,K}\leq\frac{h_{i}^{K}}{d+2}\quad\text{ for all }K\in\mathcal{T}_{h},$ (94)
@@
For $i\neq j$, from (26) and (21)
$M_{ij}=\sum_{K\in\omega_{i}\cap\omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right).$ (95)
Applying (49), we have
$M_{ij}\geq\sum_{K\in\omega_{i}\cap\omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}-\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\,\|b\|_{\infty,K}\right),$ (96)
which is nonnegative under condition (94). ∎
Combining the previous two lemmas, we have the following result immediately.
@@
Assume the hypotheses of Lemmas 4.5 and 4.6. Then
$M_{ij}-\Delta t(1-\theta)A_{ij}\geq 0,\quad i\neq j.$ (97)
@@ 4.3 Temporal estimates
Define
$\eta_{ij}^{K}:=\frac{|K|}{h_{i}^{k}h_{j}^{K}}\delta_{ij}^{K},\quad\mu_{ij}^{K}:=\frac{|K|}{(d+1)(d+2)}+\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\|b\|_{\infty,K}.$ (98)
@@
Assume that for each element $K$ in $\mathcal{T}_{h}$:
  • (i)
    The elementwise tensors $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ are symmetric and positive definite.
  • (ii)
    The meshes associated with $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy the anisotropic nonobtuse angle condition (ANAC).
  • (iii)
    The stabilization parameter $\tau_{K}$ satisfies
    $\tau_{K}\|b\|_{\infty,K}\leq C_{1}h_{K},$ (99)
    and
    $\varepsilon(\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K}))>C_{2}\frac{h^{2}_{K}}{\tau_{K}},$ (100)
    for positive constants $C_{1}$, and $C_{2}$ independent of the mesh.
  • (iv)
    For all $i\neq j$, $\delta_{ij}^{K}$ defined in (89) is positive.
Then, if
$\Delta t\geq\frac{1}{\theta}\max_{i\neq j}\frac{\sum_{K\in\omega_{i}\cap\omega_{j}}\mu_{ij}^{K}}{\sum_{K\in\omega_{i}\cap\omega_{j}}\eta_{ij}^{K}},$ (101)
the matrix $\tilde{\boldsymbol{A}}={\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}}$ is an $M$-matrix.
@@
We first show that $\tilde{\boldsymbol{A}}$ is a $Z$-matrix. For $i\neq j$, from (26) and (21)